OpenStudy (anonymous):

What is the 7th term of the geometric sequence where a1 = 1024 and a3 = 64? a. 0.015625 b. 0.0625 c. 0.25 d. 0.1 I think its c, am i right?

OpenStudy (anonymous):

Yup.

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

The constant factor between consecutive elements of the sequence is 2^-2. \[a_3=64=2^6~so~ a_7=(2^{-2})^4a_3=2^{-2}=\frac{1}{4}=0.25\]